[q20-22]

Required information

[The following information applies to the questions displayed below.]

Seiko’s current salary is $85,000. Her marginal tax rate is 32 percent and she fancies European sports cars. She purchases a new auto each year. Seiko is currently a manager for Idaho Office Supply. Her friend, knowing of her interest in sports cars, tells her about a manager position at the local BMW and Porsche dealer. The new position pays only $75,000 per year, but it allows employees to purchase one new car per year at a discount of $15,000. This discount qualifies as a nontaxable fringe benefit. In an effort to keep Seiko as an employee, Idaho Office Supply offers her a $10,000 raise. Answer the following questions about this analysis.

a. What is the annual after-tax cost to Idaho Office Supply if it provides Seiko with the $10,000 increase in salary? (Ignore payroll taxes.)

b. By how much is the offer better for Seiko on an after tax basis (Assume that Seiko is going to purchase the new car whether she switches jobs or not.)

c. What salary would Seiko need to receive from Idaho Office Supply to make her financially indifferent (after taxes) between receiving additional salary from Idaho Office Supply and accepting a position at the auto dealership? (Round your intermediate computations to the nearest dollar amount.)

After-tax cost 7,900
Offer is better by ???
Salary needed ???
Answer #1

Part B

Salary with $10,000 raise

95000

(1 – 0.32)

After­tax benefit from salary

64600

Salary

75000

(1 – 0.32)

After­tax benefit

51000

After­tax benefit of discount

15000

After­tax value of second package

66000

Offer is better by $1400 (66000-64600)

Part C

Salary needed = 95000+(1400/(1-32%)) = $95000+2059 = $97059

after­tax difference between the two offers = $1400

Salary increase

2059

× (1 – 0.32)

1400

Salary (with $10,000 + 2059 raise)

$97059

× (1 – 0.32)

After­tax benefit from salary

$66000